4(1+sin(x)=cos^2(x)
1
Expert's answer
2012-12-20T08:22:03-0500
4(1+sin(x))=cos^2(x)
we know that
cos^2(x)=1-sin^2(x)
so
4+4sin(x)=1-sin^2(x)
let sin(x)=t
3+4t+t^2=0
t=-1 or t=-3
but | sin(x) |<=1 for every x from R
sin(x)=-1
x=pi+2*pi*n, where n from Z